OpenStudy (anonymous):

What is the exact value of the expressions the square root of 180. - the square root 125. + the square root of 5.? Simplify if possible.

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

u r the smartest person online rt now can u help

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

ill take any 1s help tho

OpenStudy (anonymous):

For the first one, you can factor 180 to be a perfect square times a number and then you can pull out the perfect square. 36*5=180

OpenStudy (anonymous):

\[\sqrt{180}=\sqrt{36}*\sqrt{5}\]

OpenStudy (anonymous):

What is the exact value of the expressions the square root of 180. - the square root 125. + the square root of 5.? Simplify if possible. 2the square root of 2. 12the square root of 2. 2the square root of 5. 12the square root of 5.

OpenStudy (anonymous):

its multiple choice

OpenStudy (anonymous):

There are three different problems and only one set of multiple choice answers.

OpenStudy (anonymous):

Oh. You need to add and subtract

OpenStudy (anonymous):

Okay you just need to change all of them so that they contain the square root of 5. \[6\sqrt{5}-5\sqrt{5}+\sqrt{5}=2\sqrt{5}\]